A local finance company quotes an interest rate of 25 percent on one-year loans. So, if you borrow $30,000, the interest for the year will be $7,500. Because you must repay a total of $37,500 in one year, the finance company requires you to pay $37,500/12, or $3,125 per month over the next 12 months. Is the interest rate on this loan 25 percent? What rate would legally have to be quoted? What is the effective annual rate?

Show steps for full credit.
Answer #1

No the interest rate is not 25% it is higher
Using financial calculator
N=12
PMT=3125
PV=-30000
FV=0
CPT I/Y=3.61191%
Effective annual rate=(1+3.61191%)^12-1=53.07919%

Legally quoted rate=APR=3.61191%*12=43.34292%